A spring with mass of 2 kg has natural length of 0.5 m. A force of 25.6 N is required to maintain, it stretched to a length of 0.70 m. If the string is stretching to a length of 0.70 m and then released with initial velocity 0, find the position of mass at anytime t.

1
Expert's answer
2022-03-29T05:33:54-0400

"k=\frac Fl=51.2~\frac Nm,"

"B=l'-l=0.2~m,"

"\omega=\sqrt{\frac km}=5 .1~\frac{rad}s,"

"x=0.5+0.2\cos 5.1t."